# Navigating the Reality Gap: Privacy-Preserving On-Device Continual Adaptation of ASR for Clinical Telephony

### v3

@@ Abstract
Darshil Chauhan1, Adityasinh Solanki1, Vansh Patel1, Kanav Kapoor1, Ritvik Jain2, Aditya Bansal2, Pratik Narang1, Dhruv Kumar1 1BITS Pilani, Pilani Campus, India 2Qure.ai, India f20230609@pilani.bits-pilani.ac.in
@@ 1 Introduction
The recent surge in Self-Supervised Learning (SSL) has propelled Automatic Speech Recognition (ASR) to near-human performance on standardized benchmarks. Foundational models like Meta’s Wav2Vec 2.0 baevski2020wav2vec and OpenAI’s Whisper radford2023robust promise a future where automated transcription can digitize patient reports, allowing clinics to improve throughput with reduced operational costs. However, for specialized, high-impact clinical domains such as rural health helplines and remote clinical services, this promise remains unfulfilled. The "reality gap," representing the disparity between clean training corpora and the chaotic, noisy, and privacy-constrained environments of real-world clinical telephony, renders these state-of-the-art models practically unusable.
@@
Our baseline analysis reveals that even robust multilingual models like IndicWav2Vec javed2022indicwav2vec degrade to a prohibitively high 40.94% Word Error Rate (WER) when exposed to real-world clinical telephony audio from rural India gram_vaani. This failure is compounded by a technical deadlock: patient data privacy laws preclude the use of cloud-based adaptation services nawaz2019privacy, while rural infrastructure lacks the high-end compute required for traditional model retraining. This creates a scenario where models cannot improve because data cannot leave the local environment, and localized compute is too constrained for standard fine-tuning.
@@
To break this deadlock, we propose an on-device adaptation framework for localized, stream-based learning. We focus on data residency: ensuring that raw patient audio never leaves the local device. By leveraging Low-Rank Adaptation (LoRA) hu2022lora, we enable the model to fine-tune on incoming clinical data streams using a fraction of the trainable parameters. However, sequential adaptation in long-running clinical deployments introduces the risk of catastrophic forgetting, where the model loses its foundational linguistic capabilities mccloskey1989catastrophic. While recent benchmarks on real-world Indian speech, such as NIRANTAR javed2025nirantar, indicate that standard regularization methods (EWC, MAS) often fail to prevent this forgetting in diverse domains, we hypothesize that this instability is exacerbated by the acoustic mismatch inherent in low-bandwidth clinical telephony. To counteract this, we integrate a multi-domain experience replay mechanism chaudhry2019tiny, interleaving small buffers of general-domain data with the incoming clinical stream to anchor the learning trajectory.
@@
Furthermore, we investigate parameter-regularization strategies to stabilize this pipeline. While standard quadratic formulations (EWC) can be sensitive to the high-variance gradients of noisy telephony speech, we observe that a balanced importance estimation (Absolute Fisher) hsu2023absolute provides a more resilient regularization signal for this localized application. Our results demonstrate that these strategies yield a 17.1% relative reduction in WER on the target clinical domain, effectively stabilizing the model’s performance while respecting rigid data residency constraints. The primary research contributions of our work are:
@@ 2.1 Self-Supervised Acoustic Foundations and the Domain Gap
Low-resource speech recognition increasingly relies on Self-Supervised Learning (SSL) to leverage unlabeled audio. We build upon Wav2Vec 2.0 (baevski2020wav2vec), which achieves high data efficiency through contrastive learning; baevski2020wav2vec showed that fine-tuning on just ten minutes of data yields competitive performance on benchmarks like Librispeech (panayotov2015librispeech). To support linguistic diversity, we utilize IndicWav2Vec (javed2022indicwav2vec), scaling the architecture to cover 40 Indian languages.
@@
Despite these capabilities, a “usability gap” exists between laboratory benchmarks and real-world deployment. Clinical ASR errors are particularly risky because they affect medically salient information such as dosages, negations, and symptom reporting finley2018automatic; shafran2003challenges. This is exacerbated in rural healthcare by noisy telephonic audio and diverse regional dialects. Recent large-scale efforts such as NIRANTAR javed2025nirantar have begun characterizing these challenges across 22 Indian languages. We utilize the Gram Vaani ASR dataset (bhanushali2022gramvaani) to represent these challenges; bhanushali2022gramvaani highlight that standard models struggle to generalize to such 8kHz telephony audio without targeted adaptation.
While signal-level enhancement strategies, such as bandwidth expansion and speech super-resolution (lin2023noise; li2019audio), offer an alternative by reconstructing high-frequency components, they introduce significant computational overhead. Deploying a separate BWE model (often GAN or diffusion-based) alongside the ASR system contradicts the strict low-latency requirements of edge devices. Furthermore, privacy constraints preclude cloud-based services (nawaz2019privacy), creating a deadlock where models cannot improve because data cannot leave the local environment.
@@ 2.2 Parameter-Efficient Fine-Tuning (PEFT) and the Efficiency Gap
To address this, we adopt Low-Rank Adaptation (LoRA) (hu2022lora), which enables local fine-tuning on modest clinical hardware. This approach is essential for meeting healthcare Data Residency requirements, as it allows for localized, high-performance adaptation without needing the centralized data storage often required by federated or cloud-based approaches rieke2020future. Recent work by song2024lora on LoRA-Whisper validates this approach for multilingual ASR, demonstrating that PEFT can effectively adapt large Transformers without catastrophic interference. By constraining optimization to a low-dimensional subspace, LoRA resolves deployment challenges in privacy-sensitive clinics.
While PEFT solves efficiency, self-improving systems face the stability gap of Continual Learning (CL), specifically Catastrophic Forgetting, where new training erodes previously learned capabilities (mccloskey1989catastrophic). In healthcare, ensuring model stability over time is a critical safety requirement for clinical evaluation cycles delange2021continual. In ASR, naive updates on recent telephony transcripts lead to overfitting on specific speakers or acoustic conditions. Benchmarks like NIRANTAR javed2025nirantar further underscore this, demonstrating that no single CL strategy currently yields consistent performance across the diverse shifts encountered in real-world Indian speech.
To mitigate this, we employ a hybrid continual learning strategy that synergizes data-level and parameter-level consolidation. First, to address the distribution shift between segments, we implement Experience Replay (ER) (chaudhry2019efficient). Building on recent regularized CL studies for Indic ASR (adethya2025study) and episodic memory frameworks (yang2022online), we utilize a multi-domain experience replay mechanism. While yang2022online focused on gradient projection, we find that a direct data mixing strategy effectively mitigates catastrophic forgetting when combined with PEFT. Our buffer retains a mixture of “hard” examples (high-loss) from the target domain (Gram Vaani) and gender-balanced samples from a clean auxiliary domain (Kathbath).
Complementing this, we constrain parameter drift using Online LoRA-EWC (wang2023language), utilizing the Absolute Fisher importance estimation (hsu2023absolute). We estimate importance using the absolute value of accumulated gradients ($|g|$) rather than the standard squared Fisher Information. While this linear scaling draws inspiration from the numerical stability of Memory Aware Synapses (MAS) (aljundi2018memory), Absolute Fisher explicitly ties the importance to the likelihood of the adaptation task rather than unsupervised output sensitivity. By combining this regularized signal with multi-domain replay, our approach prevents catastrophic forgetting through a dual mechanism: replay maintains the data distribution, while the regularization penalty preserves critical adapter parameters.
@@ 3.3 Parameter-Driven Stability: Elastic Weight Consolidation (EWC)
We compute the importance of each LoRA parameter $\theta_{i}$ using the diagonal of the Fisher Information Matrix ($F$). To ensure robust optimization against high-variance gradients, we approximate importance using accumulated absolute gradients rather than squared gradients. This “Absolute Fisher” approach (hsu2023absolute) functions as a reliability optimization. We distinguish this from Memory Aware Synapses (MAS) (aljundi2018memory); while MAS assesses importance based on the sensitivity of the output function, Absolute Fisher derives importance from a linearized measure of the likelihood’s Fisher Information, preventing transient acoustic artifacts from drowning out meaningful weights:
@@ 4.1.1 Metrics
$\text{WER}=\frac{S_{w}+D_{w}+I_{w}}{N_{w}},\quad\text{CER}=\frac{S_{c}+D_{c}+I_{c}}{N_{c}}$
@@ 4.1.2 Datasets
  • Gram Vaani (Clinical Helpline Proxy) (gram_vaani): This dataset consists of rural telephonic speech (originally 8kHz, upsampled to 16kHz) and serves as a proxy for audio encountered in frontline healthcare helplines. We evaluate on the official 3-hour Evaluation set (GV_Eval_3h). Note that prior work (patel2022gramvaani) reported 30.3% WER on the 5-hour Development set using full Conformer models. Our result (33.94%) reflects performance on the unseen Evaluation partition under strictly constrained on-device conditions (LoRA vs. full tuning). To strictly simulate a continual clinical learning scenario, we partition the 103 hours of training data into sequential segments, processing them one by one to mimic a live data stream while ensuring patient data residency.
  • Kathbath (General Domain): A high-quality, read speech dataset representing standard Hindi (kathbath_hindi). We utilize a subset of the training set (approx. 25,800 samples) to populate the experience replay buffer, ensuring the model retains knowledge of standard Hindi. The complete validation set (3,151 samples) is used exclusively to measure catastrophic forgetting after adaptation.
@@ 4.2.3 Paradigm 3: Elastic Weight Consolidation (EWC)
We examine whether parameter-regularization can implicitly protect the model’s foundational knowledge without replaying data. We utilize the Absolute Fisher importance ($F_{i}$) (hsu2023absolute) to penalize changes to weights deemed critical for the source domain.
@@ 5.6 Clinical Implications and Deployment Feasibility
The 17.1% relative improvement in WER demonstrated by our Multi-Domain ER strategy (V3.1) represents a significant bridge between "raw" and "usable" clinical ASR. In a clinical dictation setting, this reduction minimizes the manual correction burden on healthcare workers, whose time is a critical resource in rural India. By enabling the system to stabilize on local dialects and noisy line conditions without transmitting patient data to central servers, we provide a technical blueprint for scalable, privacy-compliant medical transcription. We note that prior work on the Gram Vaani dataset patel2022gramvaani achieved a lower WER (30.3%) using heavy-duty architectures (Conformer/TDNN) and external Language Models. However, our framework (33.94%) operates under strict on-device constraints without external LMs during training, reducing computational footprint while ensuring adherence to rigid data residency requirements. Furthermore, the ability to run these adaptations on mobile GPUs ensures that hospitals with limited digital infrastructure can still leverage state-of-the-art ASR foundational models.
@@ 7 Privacy, Ethics, and Clinical Safety Considerations
Our primary focus on on-device adaptation is a direct response to healthcare data residency requirements. By ensuring that raw patient audio never leaves the local environment, we mitigate the risk of centralized data breaches and maintain compliance with emerging data protection laws in various jurisdictions (nawaz2019privacy; rieke2020future).
@@ A.1 Hyperparameter Configurations
Table 3 details the parameters used across our investigation. We utilize a split-table format to distinguish between universal settings and those tailored to specific continual learning strategies.

### v4

@@ Abstract
Darshil Chauhan1, Adityasinh Solanki1, Vansh Patel1, Kanav Kapoor1 Ritvik Jain2, Aditya Bansal2, Pratik Narang1, Dhruv Kumar1 1BITS Pilani, Pilani Campus, India 2Qure.ai, India f20230609@pilani.bits-pilani.ac.in
@@ 1 Introduction
The recent surge in Self-Supervised Learning (SSL) has propelled Automatic Speech Recognition (ASR) to near-human performance on standardized benchmarks. Foundational models like Meta’s Wav2Vec 2.0 Baevski et al. (2020) and OpenAI’s Whisper Radford et al. (2023) promise a future where automated transcription can digitize patient reports, allowing clinics to improve throughput with reduced operational costs. However, for specialized, high-impact clinical domains such as rural health helplines and remote clinical services, this promise remains unfulfilled. The "reality gap," representing the disparity between clean training corpora and the chaotic, noisy, and privacy-constrained environments of real-world clinical telephony, renders these state-of-the-art models practically unusable.
@@
Our baseline analysis reveals that even robust multilingual models like IndicWav2Vec Javed et al. (2022) degrade to a prohibitively high 40.94% Word Error Rate (WER) when exposed to real-world clinical telephony audio from rural India Bhanushali et al. (2022). This failure is compounded by a technical deadlock: patient data privacy laws preclude the use of cloud-based adaptation services Leroy et al. (2019), while rural infrastructure lacks the high-end compute required for traditional model retraining. This creates a scenario where models cannot improve because data cannot leave the local environment, and localized compute is too constrained for standard fine-tuning.
@@
To break this deadlock, we propose an on-device adaptation framework for localized, stream-based learning. We focus on data residency: ensuring that raw patient audio never leaves the local device. By leveraging Low-Rank Adaptation (LoRA) Hu et al. (2022), we enable the model to fine-tune on incoming clinical data streams using a fraction of the trainable parameters. However, sequential adaptation in long-running clinical deployments introduces the risk of catastrophic forgetting, where the model loses its foundational linguistic capabilities McCloskey and Cohen (1989). While recent benchmarks on real-world Indian speech, such as NIRANTAR Javed et al. (2025), indicate that standard regularization methods (EWC, MAS) often fail to prevent this forgetting in diverse domains, we hypothesize that this instability is exacerbated by the acoustic mismatch inherent in low-bandwidth clinical telephony. To counteract this, we integrate a multi-domain experience replay mechanism Chaudhry et al. (2019b), interleaving small buffers of general-domain data with the incoming clinical stream to anchor the learning trajectory.
@@
Furthermore, we investigate parameter-regularization strategies to stabilize this pipeline. While standard quadratic formulations (EWC) can be sensitive to the high-variance gradients of noisy telephony speech, we observe that a balanced importance estimation (Absolute Fisher) Benzing (2021) provides a more resilient regularization signal for this localized application. Our results demonstrate that these strategies yield a 17.1% relative reduction in WER on the target clinical domain, effectively stabilizing the model’s performance while respecting rigid data residency constraints. The primary research contributions of our work are:
@@ 2.1 Self-Supervised Acoustic Foundations and the Domain Gap
Low-resource speech recognition increasingly relies on Self-Supervised Learning (SSL) to leverage unlabeled audio. We build upon Wav2Vec 2.0 (Baevski et al., 2020), which achieves high data efficiency through contrastive learning; Baevski et al. (2020) showed that fine-tuning on just ten minutes of data yields competitive performance on benchmarks like Librispeech (Panayotov et al., 2015). To support linguistic diversity, we utilize IndicWav2Vec (Javed et al., 2022), scaling the architecture to cover 40 Indian languages.
@@
Despite these capabilities, a “usability gap” exists between laboratory benchmarks and real-world deployment. Clinical ASR errors are particularly risky because they affect medically salient information such as dosages, negations, and symptom reporting Chiu et al. (2018a, b). This is exacerbated in rural healthcare by noisy telephonic audio and diverse regional dialects. Recent large-scale efforts such as NIRANTAR Javed et al. (2025) have begun characterizing these challenges across 22 Indian languages. We utilize the Gram Vaani ASR dataset (Bhanushali et al., 2022) to represent these challenges; Bhanushali et al. (2022) highlight that standard models struggle to generalize to such 8kHz telephony audio without targeted adaptation.
While signal-level enhancement strategies, such as bandwidth expansion and speech super-resolution (Lin et al., 2023; Li et al., 2019), offer an alternative by reconstructing high-frequency components, they introduce significant computational overhead. Deploying a separate BWE model (often GAN or diffusion-based) alongside the ASR system contradicts the strict low-latency requirements of edge devices. Furthermore, privacy constraints preclude cloud-based services (Leroy et al., 2019), creating a deadlock where models cannot improve because data cannot leave the local environment.
@@ 2.2 Parameter-Efficient Fine-Tuning (PEFT) and the Efficiency Gap
To address this, we adopt Low-Rank Adaptation (LoRA) (Hu et al., 2022), which enables local fine-tuning on modest clinical hardware. This approach is essential for meeting healthcare Data Residency requirements, as it allows for localized, high-performance adaptation without needing the centralized data storage often required by federated or cloud-based approaches Rieke et al. (2020). Recent work by Song et al. (2024) on LoRA-Whisper validates this approach for multilingual ASR, demonstrating that PEFT can effectively adapt large Transformers without catastrophic interference. By constraining optimization to a low-dimensional subspace, LoRA resolves deployment challenges in privacy-sensitive clinics.
While PEFT solves efficiency, self-improving systems face the stability gap of Continual Learning (CL), specifically Catastrophic Forgetting, where new training erodes previously learned capabilities (McCloskey and Cohen, 1989). In healthcare, ensuring model stability over time is a critical safety requirement for clinical evaluation cycles De Lange et al. (2022). In ASR, naive updates on recent telephony transcripts lead to overfitting on specific speakers or acoustic conditions. Benchmarks like NIRANTAR Javed et al. (2025) further underscore this, demonstrating that no single CL strategy currently yields consistent performance across the diverse shifts encountered in real-world Indian speech.
To mitigate this, we employ a hybrid continual learning strategy that synergizes data-level and parameter-level consolidation. First, to address the distribution shift between segments, we implement Experience Replay (ER) (Chaudhry et al., 2019a). Building on recent regularized CL studies for Indic ASR (T and Nirmala, 2025) and episodic memory frameworks (Yang et al., 2022), we utilize a multi-domain experience replay mechanism. While Yang et al. (2022) focused on gradient projection, we find that a direct data mixing strategy effectively mitigates catastrophic forgetting when combined with PEFT. Our buffer retains a mixture of “hard” examples (high-loss) from the target domain (Gram Vaani) and gender-balanced samples from a clean auxiliary domain (Kathbath).
Complementing this, we constrain parameter drift using Online LoRA-EWC (Xiang et al., 2023), utilizing the Absolute Fisher importance estimation (Benzing, 2021). We estimate importance using the absolute value of accumulated gradients ($|g|$) rather than the standard squared Fisher Information. While this linear scaling draws inspiration from the numerical stability of Memory Aware Synapses (MAS) (Aljundi et al., 2018), Absolute Fisher explicitly ties the importance to the likelihood of the adaptation task rather than unsupervised output sensitivity. By combining this regularized signal with multi-domain replay, our approach prevents catastrophic forgetting through a dual mechanism: replay maintains the data distribution, while the regularization penalty preserves critical adapter parameters.
@@ 3.3 Parameter-Driven Stability: Elastic Weight Consolidation (EWC)
We compute the importance of each LoRA parameter $\theta_{i}$ using the diagonal of the Fisher Information Matrix ($F$). To ensure robust optimization against high-variance gradients, we approximate importance using accumulated absolute gradients rather than squared gradients. This “Absolute Fisher” approach (Benzing, 2021) functions as a reliability optimization. We distinguish this from Memory Aware Synapses (MAS) (Aljundi et al., 2018); while MAS assesses importance based on the sensitivity of the output function, Absolute Fisher derives importance from a linearized measure of the likelihood’s Fisher Information, preventing transient acoustic artifacts from drowning out meaningful weights:
@@ 4.1.1 Metrics
$\text{WER}=\frac{S_{w}+D_{w}+I_{w}}{N_{w}}$
$\text{CER}=\frac{S_{c}+D_{c}+I_{c}}{N_{c}}$
@@ 4.1.2 Datasets
  • Gram Vaani (Clinical Helpline Proxy) (Bhanushali et al., 2022): This dataset consists of rural telephonic speech (originally 8kHz, upsampled to 16kHz) and serves as a proxy for audio encountered in frontline healthcare helplines. We evaluate on the official 3-hour Evaluation set (GV_Eval_3h). Note that prior work (Patel and Scharenborg, 2022) reported 30.3% WER on the 5-hour Development set using full Conformer models. Our result (33.94%) reflects performance on the unseen Evaluation partition under strictly constrained on-device conditions (LoRA vs. full tuning). To strictly simulate a continual clinical learning scenario, we partition the 103 hours of training data into sequential segments, processing them one by one to mimic a live data stream while ensuring patient data residency.
  • Kathbath (General Domain): A high-quality, read speech dataset representing standard Hindi (Javed et al., 2023). We utilize a subset of the training set (approx. 25,800 samples) to populate the experience replay buffer, ensuring the model retains knowledge of standard Hindi. The complete validation set (3,151 samples) is used exclusively to measure catastrophic forgetting after adaptation.
@@ 4.2.3 Paradigm 3: Elastic Weight Consolidation (EWC)
We examine whether parameter-regularization can implicitly protect the model’s foundational knowledge without replaying data. We utilize the Absolute Fisher importance ($F_{i}$) (Benzing, 2021) to penalize changes to weights deemed critical for the source domain.
@@ 5.6 Clinical Implications and Deployment Feasibility
The 17.1% relative improvement in WER demonstrated by our Multi-Domain ER strategy (V3.1) represents a significant bridge between "raw" and "usable" clinical ASR. In a clinical dictation setting, this reduction minimizes the manual correction burden on healthcare workers, whose time is a critical resource in rural India. By enabling the system to stabilize on local dialects and noisy line conditions without transmitting patient data to central servers, we provide a technical blueprint for scalable, privacy-compliant medical transcription. We note that prior work on the Gram Vaani dataset Patel and Scharenborg (2022) achieved a lower WER (30.3%) using heavy-duty architectures (Conformer/TDNN) and external Language Models. However, our framework (33.94%) operates under strict on-device constraints without external LMs during training, reducing computational footprint while ensuring adherence to rigid data residency requirements. Furthermore, the ability to run these adaptations on mobile GPUs ensures that hospitals with limited digital infrastructure can still leverage state-of-the-art ASR foundational models.
@@ 7 Privacy, Ethics, and Clinical Safety Considerations
Our primary focus on on-device adaptation is a direct response to healthcare data residency requirements. By ensuring that raw patient audio never leaves the local environment, we mitigate the risk of centralized data breaches and maintain compliance with emerging data protection laws in various jurisdictions (Leroy et al., 2019).
@@ A.1 Hyperparameter Configurations
Table 4 details the parameters used, utilizing a split-table format to distinguish between universal settings and those tailored to specific continual learning strategies.
A.2 Algorithm and Buffer Management
The technical contribution of our framework lies in the efficient integration of LoRA, linearized EWC, and prioritized experience replay. Algorithm 1 below details the localized adaptation loop, and Algorithm 2 describes our multi-domain buffer management strategy.
Algorithm 1: LoRA-based Hybrid Adaptation Loop
Require: Base model $\theta_{base}$, Replay buffer $\mathcal{B}$, Regularization $\lambda$ 1: Initialize LoRA adapters $\theta_{0}\subset\theta_{base}$ and importance $F\leftarrow\mathbf{0}$ 2: for segment $k=1,\dots,K$ do
3:
 Receive clinical stream $\mathcal{D}_{k}$
[…]
  $\theta_{k}\leftarrow\text{argmin}_{\theta}\mathcal{L}_{CTC}(\mathcal{D}_{train})$
$\frac{\lambda}{2}\sum F_{i}(\theta_{k}-\theta_{k-1}^{*})^{2}$ 6:Importance Estimation:
  $F_{new}\leftarrow\frac{1}{N}\sum_{x\in\mathcal{D}_{train}}|\nabla_{\theta}\mathcal{L}_{CTC}(x)|$
7:Asynchronous Consolidation:
  $F\leftarrow\frac{F\times(k-1)+F_{new}}{k}$
[…]
10: end for
Algorithm 2: Buffer Management
Require: Segment data $\mathcal{D}_{k}$, Buffer $\mathcal{B}$, Threshold $\tau$ 1: Compute $\mathcal{L}_{inst}(x)$ for all instances $x\in\mathcal{D}_{k}$ 2: Hard Example Mining:
3: Buffer Update:
$\mathcal{B}_{gv}\leftarrow\text{Sample}(60\%\text{ from }\mathcal{S}_{hard}$
$\cup\ 40\%\text{ Random})$ $\mathcal{B}_{gen}\leftarrow\text{Sample}(300\text{ Balanced from Kathbath})$ 4: $\mathcal{B}\leftarrow\mathcal{B}_{gv}\cup\mathcal{B}_{gen}$
A.3 Computational Efficiency and Hardware Setup
A core goal of our work is to prove that high-performance, privacy-preserving adaptation is possible on standard mobile workstations. We evaluated our pipeline on the configurations detailed in Table 5. These benchmarks demonstrate the technical feasibility of localized, self-improving ASR systems in environments where high-end compute clusters are unavailable.
Appendix B Detailed Experimental Results
Gradient Stability: A critical challenge encountered during the development of the EWC-based paradigms was gradient explosion. As shown later in Figure 13, our proposed Linearized EWC (L-EWC) strategy successfully stabilizes the gradient norm throughout the 24 adaptation segments compared to standard quadratic formulations.
Forgetting Analysis: The preservation of foundational general-domain knowledge (Kathbath) is evaluated across all experimental paradigms, with detailed comparisons provided in Figures 9 through 12.
Appendix C Dataset Characteristics
The Gram Vaani dataset serves as a rigorous proxy for rural clinical environments due to its telephonic acquisition (originally 8kHz upsampled to 16kHz) and focus on medical/agricultural discussions. Table 3 summarizes the key characteristics.
Metric Value
Total Duration 103.2 Hours
Number of Segments ($k$) 24
Samples per Segment $\approx$ 1,600
Avg. Duration per Sample 9.4 s
Acoustic Condition Noisy
Primary Dialect Rural Hindi